How many ounces of a 15% alcohol solution must be mixed with 4 ounces of a 20% alcohol solution to make a 17% alcohol solution
Answer by Earlsdon(6294)

Let x = the number of ounces of the 15% alcohol solution needed to be mixed with the 4 ounces of 20% alcohol solution to obtain the 17% alcohol solution.
After converting the percentages to their decimal equivalents, set up the necessary equation as follows:
0.15x+0.20(4) = (x+4)(0.17) Simplify and solve for x.
0.15x+0.8 = 0.17x+0.68 Subtract 0.15x from both sides.
0.8 = 0.02x+0.68 Now subtract 0.68 from both sides.
0.12 = 0.02x Finally, divide both sides by 0.02
x = 6
You need to mix 6 ounces of 15% alcohol solution with 4 ounces of 20% alcohol solution to obtain 10 (x+4)ounces of 17% alcohol solution.